OpenStudy (anonymous):

Determine the value of y for the system of equations. x − 3y = 4 2x + y = 8

hartnn (hartnn):

Multiply 1st equation by 2, what you get ?

OpenStudy (evonhowell):

x - 3y = 4 2x + y = 8 Times the first equation by -2 -2(x - 3y = 4); simplify -2x + 6y = -8 Put the two new equations together -2x + 6y = -8 2x + y = 8 (+) _________ 7y = 0; divide each term by 7 7y/7 = 0/7; solve y = 0 <-----------Answer

OpenStudy (evonhowell):

Thats what I got

OpenStudy (anonymous):

thanks(:

OpenStudy (evonhowell):

Your Welcome :)